John is currently 25 years old. He has $10,000 saved up and wishes to deposit this into a savings account which pays him J12 = 6% p.a. He also wishes to deposit $X every month into that account so that when he retires at 55, he can withdraw $2000 every month end to support his retirement. He expects to live up till 70 years. How much should he deposit every month into his account
1
Expert's answer
2020-06-04T17:16:25-0400

Amount John want to have at age 55,


"2,000*12*15=360,000"


currently he wants to invest 10,000

find the future value of the 10,000


"future \, value=present \,value * (1+r )^n"

"future \, value=10,000 * (1+0.05)^{360}"

"future \, value=60,225,"


"360,000-60,225=299,775"

John's monthly savings would total to 299,775 at the age of 55


"future \, value \, of \, annuity= present \, value *\frac{1-(1+r )^{-n}}{r}"

"299,775 = x *\frac{1-(1+0.005 )^{-360}}{0.005}"


"x=montly \, deposit"


"John's \, monthly \, deposit=1,797"